Consider a buffer prepared by mixing 75ml of 0.03640 M ammonia (NH3) solution with 28ml of 0.02454 M ammonium chloride solution (pKb of ammonia is 4.75)

1. What will the pH be?

2. If you add 1 mL of 0.1780 M HCl, what will the new pH be?
Answer #1

Total volume of the solution = (75+28) ml = 103 ml 75 ml 0,0364 M NH3 solution is mixed with 28 me 0.02454M Let us suppose NHor pot = 4.75 + log/0.0066 0.0265 POH = 4.75 + (-0.6037) 08 or poH = 4.146 We know that, pH + pot = 14 So, pH = 14- pOH = (14Here amount of NHA ce increased by 0.000178 mal and amount of NH₃ decreased by 0.000178 mol 103 ml 0.02.65 M NH = 103 x 0.026Concentration of fall of Nigel = (0.0008578x1000) So, (sall] = 0.008248 (M) 104 Sac So, on poH = PKB+ log [ sall) [base] poH